# PolicyPad: Collaborative Prototyping of LLM Policies

### v1

@@ Abstract.
As LLMs gain adoption in high-stakes domains like mental health, domain experts are increasingly consulted to provide input into policies governing their behavior. From an observation of 19 policymaking workshops with 9 experts over 15 weeks, we identified opportunities to better support rapid experimentation, feedback, and iteration for collaborative policy design processes. We present PolicyPad, an interactive system that facilitates the emerging practice of LLM policy prototyping by drawing from established UX prototyping practices, including heuristic evaluation and storyboarding. Using PolicyPad, policy designers can collaborate on drafting a policy in real time while independently testing policy-informed model behavior with usage scenarios. We evaluate PolicyPad through workshops with 8 groups of 22 domain experts in mental health and law, finding that PolicyPad enhanced collaborative dynamics during policy design, enabled tight feedback loops, and led to novel policy contributions. Overall, our work paves participatory paths for advancing AI alignment and safety.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Synchronous editors
Figure 1. PolicyPad is an interactive system that facilitates collaborative prototyping of LLM policies. Policy designers work together in real time (left) to draft policy statements in PolicyPad’s collaborative editor (middle), while experimenting with the model’s policy-informed behavior in a private sidebar (right). Content from the private sidebar can be fluidly brought into the collaborative editor for viewing, editing, and discussion. To facilitate LLM policy prototyping, PolicyPad borrows concepts and practices from UX prototyping, including heuristic evaluation, storyboarding, and rapid iteration.:
@@ 1. Introduction
For decades, researchers and science fiction writers have imagined a world in which AI systems can be governed by natural language rules (Asimov, 1940; Turing, 1950). Today, governing large language models (LLMs) with LLM policies—sets of rules, guidelines, and desiderata that shape model behavior—is a key component in the broader toolkit of approaches to improve model alignment and safety (OpenAI, 2025, 2025a; Anthropic, 2023; Huang et al., 2024; Horwitz, 2025; Lam et al., 2024). For example, OpenAI’s Model Spec contains a series of general objectives and principles (e.g., “Seek the truth together”) for researchers and red-teamers to use as a guide when working on reinforcement learning from human feedback (RLHF) (OpenAI, 2025a), as well as for the model to learn from directly (Guan et al., 2024). Similarly, Anthropic uses Constitutional AI—in which reinforcement learning receives reward signals from AI-generated feedback that adheres to a set of principles (a “constitution”)—to align its Claude models (Bai et al., 2022; Anthropic, 2023). If effective, LLM policies promise a transparent, familiar, and legible means by which developers and policymakers can govern AI systems.
@@
As LLMs are deployed to millions of users globally, LLM policies become increasingly consequential and scrutinized. This is especially true in high-stakes, tightly regulated domains that are seeing rapid increases in LLM use by everyday users, such as mental health and law (Hill, 2025a; Lohr, 2023; Rosenbluth, 2025; Hill and Freedman, 2025; Hill, 2025b; Cheong et al., 2024; Perez, 2025). LLM policies are primarily written by model developers, but the lack of external input often leads to insular policies that poorly reflect the attitudes and values of everyday users, while ignoring key safety concerns (Huang et al., 2024; Horwitz, 2025). While frontier AI labs regularly partner with external domain experts to conduct pre-release safety testing of their models (Anthropic, 2025b; Hurst et al., 2024; Google, 2025), there has been little documentation of similar efforts for LLM policies. Yet, there is mounting recognition that expert input is essential for LLM policy design, especially in safety-critical, domain-specific use cases (OpenAI, 2025b; Walsh and Gronvall, 2025).
@@
In this work, we first conduct a 15-week observational study in partnership with a U.S.-based frontier AI lab to better understand how domain experts can contribute to LLM policy design. Through 19 interactive workshops in which mental health experts discussed, annotated, taxonomized, and drafted user queries and LLM responses, we observed experts collaboratively ideating and discussing policy ideas while seeking ways to rapidly test and iterate on them through experimentation with model behavior on realistic scenarios. Much like prototyping in user experience (UX) practice, there is a strong emphasis on collaborative and rapid exploration, feedback collection, and iteration. We thus conceptualize this emerging practice as LLM policy prototyping, borrowing from established UX practices like heuristic evaluation and low-fidelity prototyping.
@@
However, few tools exist for policy design (Lam et al., 2024), let alone tools that support collaborative LLM policy prototyping. We observe notable opportunities for experts to tighten the feedback loop during policy design while leveraging collaborative affordances to build off each other’s expertise. We design and develop PolicyPad, [^1] We plan to open-source PolicyPad at https://github.com/kjfeng/policypad. an interactive system that facilitates LLM policy prototyping. PolicyPad draws upon established methods and concepts within UX prototyping to enable small groups to collaboratively draft policies, test policy-informed model behavior against usage scenarios, evaluate the quality of the policy, and iterate on its contents in real time through tight feedback loops.
@@
We evaluate PolicyPad through policy prototyping sessions with 22 domain experts spanning two domains—mental health and law—organized into 8 groups. We found that design decisions in PolicyPad fostered collaborative dynamics between experts via its interactive in-editor widgets and yielded more novel policies compared to a baseline, relative to existing policies [^2] As of September 1, 2025. including OpenAI’s Model Spec (OpenAI, 2025a) and Claude’s Constitution (Anthropic, 2023). Key areas of novelty include offering more specific guidelines on when the model should defer to a human expert, and eliciting key information required for responsible assistance early in the conversation. We end by discussing the practical implications of our work, including where LLM policy prototyping can be situated in AI alignment pipelines, and approaches for scaling up policy prototyping efforts.
@@
  • An 15-week observational study with 9 mental health experts that surfaced opportunities for tight, collaborative feedback loops in LLM policy design.
@@
Figure 2. Research Process Overview. Our work proceeded in 4 phases: (1) a 15-week observational study with 9 mental health experts (19 workshops) led to (2) conceptualization of LLM policy prototyping. We then (3) designed and built PolicyPad and (4) evaluated it through 8 policy prototyping sessions with 22 experts (10 mental health, 12 legal).:
@@ 2. Related Work
2.1. Participatory and Collaborative Policy Design
Policies are instruments of governance to guide decisions and establish desirable goals and outcomes (Ostrom, 1990). Research has shown that a lack of input from broader constituents has led to governance policies that are poorly informed, fail to address concerns of those affected, and lack legitimacy and trust in the public eye (Moon, 2023; Lazar, 2024; Bua and Escobar, 2018). Thus, expanding the policy-making processes to a wider range of stakeholders through the participatory design of policy can confer many benefits, such as achieving more democratic legitimacy (Ovadya et al., 2025), providing better outputs through the integration of localized or specialized knowledge (Landemore and Fourniau, 2022), and building social cohesion (Michels and Graaf, 2010). Participatory policy design can be especially crucial for public-facing AI and LLM systems (Delgado et al., 2023; Suresh et al., 2024), as these systems impact large portions of the population (Makridakis, 2017; Jošt et al., 2024; De Choudhury et al., 2023), while also having many domain-specific risks due to their broad utility (Bommasani et al., 2021; Hart, 2025; Hill, 2025a).
Participatory policy design has been practiced in offline governance, with structured processes deployed to create public policies around transportation, environment, technology, and public health around the world (People Powered, 2025; Democracy Beyond Elections, 2021). Expert workshops have been one common way to achieve participatory inputs in offline settings, with prominent applications in business (Hodgkinson et al., 2006), public health (Rushmer et al., 2014), and politics (Andersen and Jæger, 1999). Deliberative democracy takes this a step further, where representative groups of community members not only contribute knowledge but also deliberate over policy issues (Ryfe, 2005; Elster, 1998; Gutmann and Thompson, 2004). In particular, citizens’ assemblies—where small samples of the population participate in workshops where they receive briefings about an issue before deliberating on policy options—have proven to be effective in producing agreeable and actionable policy reports on complex topics while still engaging with largely laypeople constituents (Fournier, 2011; Smith and Wales, 2000; Landemore and Fourniau, 2022; Lenaghan, 1999; Public Participation Guide: Citizen Juries, 2025).
There have been many digital tools developed to augment, scale up, or improve on certain aspects of participatory policy design (Goldberg et al., 2024). Some efforts address scaling by aggregating and distilling opinions from lower-level preference inputs (Bai et al., 2022; Anthropic, 2023; OpenAI, 2023). Moving beyond simple opinion aggregation, tools like Pol.is (Small et al., 2021), Talk to the City (AI Objectives Institute, 2023), Remesh (Remesh Inc., 2025; Konya et al., 2023), and ConsiderIt (Kriplean et al., 2012) also include mechanisms for consensus-finding and balancing of trade-offs. Various tools have also been developed to assist and engage online communities in authoring policies in text or code (Zhang et al., 2020; Mahyar et al., 2018; Kuo et al., 2024) with others supporting decision-making processes with policies or policy-like guidance (Fan and Zhang, 2020; Koshy et al., 2024; Chen and Zhang, 2025).
Participatory processes with stakeholder inputs have more recently also been applied to the governance of technology and online platforms. For example, platforms like Facebook have delegated some aspects of content moderation policy to panels with external stakeholders such as the Oversight Board (Klonick, 2019; Price and Price, 2023). Other online platforms take a decentralized approach, enabling communities to create and enforce their own policies (Jhaver et al., 2023). Government adoption of AI systems has also faced considerable public deliberation and debate in the US (Washington, 2018; Arnesen et al., 2024) and EU (van Noordt and Misuraca, 2022). Increasingly, such policy processes are adopting tools and approaches from participatory policy design, including public surveys and participatory expert workshops in the UK (Marres et al., 2024) and a citizens’ assembly on AI in the EU  (Graham Wetherall-Grujić, 2024).
Current practical efforts around designing policies for governing LLM behavior have largely been directed by actors within the AI industry. In an effort to ensure the robustness of LLM policies, AI developers and operators often employ in-house teams to surface policy weaknesses through “red-teaming” efforts (Ahmad et al., 2025; Feffer et al., 2024), though many external users and testers may also voluntarily contribute to identifying and reporting such issues as well (Zhang et al., 2024). More recently, there have also been pushes to formalize safety policies around AI systems via compliance testing organizations (Team, 2021) and from the academic community (Ozlati and Yampolskiy, 2017; Anderljung et al., 2023). Organizations have also involved the general public through programs aimed at collecting public input on specific topics of policy concern (OpenAI, 2025; Anthropic, 2023; OpenAI, 2023). However, compared to offline participatory efforts, a notably missing component is better integration of domain experts—practitioners who have specialized expertise and field experience, but not technical expertise on AI. Compared to laypeople, these domain experts can often give more nuanced input for safety-critical use cases such as mental health, and are thus promising collaborators for designing LLM policies (Cheong et al., 2024). We thus contribute tooling to support this collaboration in our work.
2.2. Theory and Practice of Prototyping
Prototypes are fundamental artifacts in the design of physical and digital products. At their core, prototypes are representations of an interactive system that simulates one or more aspects of its design and functionality (Lim et al., 2008; Lauff et al., 2018; Beaudouin-Lafon and Mackay, 2002; Figma, 2025b; Interaction Design Foundation, 2025; Houde and Hill, 1997). Depending on what a designer hopes to learn at a particular point in the design process, the prototypes they employ may vary in fidelity, interactivity, duration of use, or representation (Rudd et al., 1996; Beaudouin-Lafon and Mackay, 2002; Houde and Hill, 1997). For example, prototypes may represent the look and feel of a product, the product’s role in users’ lives, implementation approaches, or some combination thereof (Houde and Hill, 1997). Lim et al.posit that the most effective prototype is economical, one that “makes the possibilities and limitations of a design idea visible and measurable […] in the simplest and most efficient way”  (Lim et al., 2008).
Successful prototypes not only help individual designers explore, conceptualize, and validate design ideas, but also act as boundary objects (Star and Griesemer, 1989) that facilitate dialogue and collaboration among groups of stakeholders involved in the design process (Feng et al., 2023, 2024; Subramonyam et al., 2022; Liao et al., 2023; Park et al., 2022; Delgado et al., 2023; Buchenau and Suri, 2000). Specifically, the production of tentative designs that can be quickly iterated upon is an essential component in aligning ideas and expectations in participatory processes (Delgado et al., 2023; Mackay and Beaudouin-Lafon, 2023). The fidelity of prototypes also significantly impacts their communicative role. Low-fidelity prototypes, such as paper sketches and wireframes, support rapid and parallel ideation by teams early in the design process, while the interactivity and details in high-fidelity prototypes yield more granular feedback to help teams make their ideas more concrete (Dow et al., 2010; Rudd et al., 1996). However, researchers and practitioners have also argued against strict adherence to a low-to-high fidelity prototyping workflow. For example, Virzi et al. (Virzi et al., 1996) found that low-fidelity prototypes can provide effective feedback to designers throughout the product design lifecycle, not just in the early stages. The fidelity level at which prototyping happens should thus be calibrated based on the designers’ goals and problems faced (Lim et al., 2008; Beaudouin-Lafon and Mackay, 2002; Houde and Hill, 1997).
In recent years, researchers and practitioners have started to draw from prototyping practices to improve policymaking processes (Kimbell and Bailey, 2017; Hagan, 2021). Policy prototyping has been proposed to enable policymakers to iteratively collect and incorporate feedback on early policy drafts (Project Let’s Talk Privacy, 2020; Quicksey and Meierling, 2022; Kontschieder, 2018). However, testing and evaluating prototypes of traditional policies often remain intractable in practice due to the need for deployment on real populations and long evaluation timelines. Given that LLM behaviors can be quickly adapted and tested within tight feedback loops for proof-of-concept experimentation and evaluation (Anthropic, 2025a; Feng et al., 2024), we observe that LLM policies are an especially promising candidate for policy prototyping.
2.3. Methods for Prompt Engineering and Red Teaming
At first glance, LLM policy design shares a key goal with the popular practice of prompt engineering: to produce a natural language artifact that shapes model behavior (Zamfirescu-Pereira et al., 2023; OpenAI, 2025a). However, the two have fundamentally different epistemic goals.
In prompt engineering, the goal is output-driven—to create a prompt that elicits outputs from a model that satisfy certain criteria (Arawjo et al., 2024; Strobelt et al., 2023; Mishra et al., 2025; Jiang et al., 2022). The design of interactive tools for prompt engineering reflects this goal. Tools like Chainforge (Arawjo et al., 2024), EvalLM (Kim et al., 2024), and CoPrompter (Joshi et al., 2025) all allow users to set customized evaluation criteria to compare and evaluate different prompt versions. Because the consumer of the final artifact is the model, factors like legibility and clarity to humans are secondary (Anthropic, 2025c); the best prompt is one that produces model outputs that best satisfy evaluation criteria (Arawjo et al., 2024; Kim et al., 2024; Joshi et al., 2025; Mishra et al., 2025).
On the other hand, the goal of LLM policy design in our work is input-driven—to elicit and encode perspectives on responsible model behavior into a concrete artifact that informs a set of practices for model alignment. The primary consumers of the policy are humans, although the policy may sometimes be used to instruct models too (Guan et al., 2024). As a result, factors like legibility and clarity are crucial. LLM policy design may benefit from some prompt engineering practices, such as immediately updating model behavior for quick testing and iteration, applying updates across a range of test cases for evaluation, and version tracking (Kim et al., 2024; Joshi et al., 2025; Arawjo et al., 2024). We thus draw design inspiration from some prompt engineering tools in our work. However, we also design for noteworthy differences, such as scaffolding evaluation of the policy rather than model outputs.
Related to prompt engineering is the practice of red teaming LLMs, where prompts—often adversarial—are perturbed and sent to the model at scale to evaluate the robustness of model safeguards (Dinan et al., 2019; Ganguli et al., 2022; Casper et al., 2023; Perez et al., 2022). Like human-centered design, red teaming often benefits from exploring diverse user scenarios to envision edge cases in model use to surface undesirable behaviors (Ganguli et al., 2022; Deng et al., 2025). However, the nature of these scenarios varies significantly across the two practices. Red teaming prioritizes scenario quantity—even minor perturbations may be useful in surfacing jailbreak vulnerabilities in models (Perez et al., 2022). On the other hand, scenarios in human-centered design serve as provocations for designers and thus prioritize quality and depth (Figma, 2025a; Carrol, 1999; Bodker, 1999; Hooper and Hsia, 1982). In our work, while some LLM policy design goals are similar to those of red teaming, we focus on supporting in-depth, small-group collaboration, with an aim to surface both undesirable and desirable model behaviors, rather than large-scale efforts specifically targeting undesirable behaviors.

### v3

@@ Abstract.
As LLMs gain adoption in high-stakes domains like mental health, domain experts are increasingly consulted to provide input into policies governing their behavior. From an observation of 19 policymaking workshops with 9 experts over 15 weeks, we identified opportunities to better support rapid experimentation, feedback, and iteration for collaborative policy design processes. We present PolicyPad, an interactive system that facilitates the emerging practice of LLM policy prototyping by drawing from established UX prototyping practices, including heuristic evaluation and storyboarding. Using PolicyPad, policy designers can collaborate on drafting a policy in real time while independently testing policy-informed model behavior with usage scenarios. We evaluate PolicyPad through workshops with 8 groups of 22 domain experts in mental health and law, finding that PolicyPad enhanced collaborative dynamics during policy design, enabled tight feedback loops, and led to novel policy contributions. Overall, our work paves expert-informed paths for advancing AI alignment and safety.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791689isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Synchronous editors
Figure 1. PolicyPad is an interactive system that facilitates collaborative prototyping of LLM policies. Policy designers work together in real time (left) to draft policy statements in PolicyPad’s collaborative editor (middle), while experimenting with the model’s policy-informed behavior in a private sidebar (right). Content from the private sidebar can be fluidly brought into the collaborative editor for viewing, editing, and discussion. To facilitate LLM policy prototyping, PolicyPad borrows concepts and practices from UX prototyping, including heuristic evaluation, storyboarding, and rapid iteration.
Policy designers work together in real time (left) to draft policy statements in PolicyPad’s collaborative editor (middle), while experimenting with the model’s policy-informed behavior in a private sidebar (right). Content from the private sidebar can be fluidly brought into the collaborative editor for viewing, editing, and discussion. To facilitate LLM policy prototyping, PolicyPad borrows concepts and practices from UX prototyping, including heuristic evaluation, storyboarding, and rapid iteration.
@@ 1. Introduction
For decades, researchers and science fiction writers have imagined a world in which AI systems can be governed by natural language rules (Asimov, 1940; Turing, 1950). Today, governing large language models (LLMs) with LLM policies [^1] In this paper, our use of the term “policy” will refer to LLM policy (as opposed to public policies, government regulations, etc.) unless stated otherwise.—sets of rules, guidelines, and desiderata that shape model behavior—is a key component in the broader toolkit of approaches to improve model alignment and safety (OpenAI, 2025a, b; Anthropic, 2023; Huang et al., 2024; Horwitz, 2025; Lam et al., 2024). For example, OpenAI’s Model Spec contains a series of general objectives and principles (e.g., “Seek the truth together”) for researchers and red-teamers to use as a guide when working on reinforcement learning from human feedback (RLHF) (OpenAI, 2025b), as well as for the model to learn from directly (Guan et al., 2024). Similarly, Anthropic uses Constitutional AI—in which reinforcement learning receives reward signals from AI-generated feedback that adheres to a set of principles (a “constitution”)—to align its Claude models (Bai et al., 2022; Anthropic, 2023). If effective, LLM policies promise a transparent, familiar, and legible means by which developers and policymakers can govern AI systems (He et al., 2025; Zhang et al., 2025b; Huang et al., 2024).
@@
As LLMs are deployed to millions of users globally, LLM policies become increasingly consequential and scrutinized. This is especially true in high-stakes, tightly regulated domains that are seeing rapid increases in LLM use by everyday users, such as mental health and law (Hill, 2025a; Lohr, 2023; Rosenbluth, 2025; Hill and Freedman, 2025; Hill, 2025b; Cheong et al., 2024; Perez, 2025). LLM policies are primarily written by model developers, but the lack of expert input often leads to insular policies that risk delivering irresponsible model outputs to users in critical scenarios, while ignoring key safety concerns (Horwitz, 2025; Hill and Freedman, 2025; Cheong et al., 2024). While frontier model developers regularly partner with external domain experts to conduct pre-release safety testing of their models (Anthropic, 2025; Hurst et al., 2024; Google, 2025), there has been little documentation of similar efforts for LLM policies. Yet, there is mounting recognition that co-designing AI behaviors with experts is essential for LLM policy design, especially in safety-critical, domain-specific use cases (OpenAI, 2025c; Walsh and Gronvall, 2025).
@@
In this work, we first conduct a 15-week observational study in partnership with OpenAI to better understand how LLM policies can be co-designed with experts. Through 19 interactive workshops in which mental health experts discussed, annotated, taxonomized, and drafted user queries and LLM responses, we observed experts collaboratively ideating and discussing policy ideas while seeking ways to rapidly test and iterate on them through experimentation with model behavior on realistic scenarios. Much like prototyping in user experience (UX) practice, there is a strong emphasis on collaborative and rapid exploration, feedback collection, and iteration. We thus conceptualize this emerging practice as LLM policy prototyping, borrowing from established UX practices like heuristic evaluation and low-fidelity prototyping. LLM policy prototyping draws upon and shares many motivations with LLM red-teaming and participatory model evaluation, but is specifically oriented towards producing an artifact that can subsequently inform red-teaming, model evaluation, and other safety efforts.
@@
However, few tools exist for policy design (Lam et al., 2024), let alone tools that support collaborative LLM policy prototyping. We observe notable opportunities for experts to tighten the feedback loop during policy design while leveraging collaborative affordances to build off each other’s expertise. We design and develop PolicyPad, [^2] We open-source PolicyPad at https://github.com/kjfeng/policypad. an interactive system that facilitates LLM policy prototyping. PolicyPad draws upon established methods and concepts within UX prototyping to enable small groups to collaboratively draft policies, test policy-informed model behavior against usage scenarios, evaluate the quality of the policy, and iterate on its contents in real time through tight feedback loops.
@@
We evaluate PolicyPad through policy prototyping sessions with 22 domain experts spanning two domains—mental health and law—organized into 8 groups. We found that design decisions in PolicyPad fostered collaborative dynamics between experts via its interactive in-editor widgets and yielded more novel policies compared to a baseline, relative to existing policies including OpenAI’s Model Spec (OpenAI, 2025b) and Claude’s Constitution (Anthropic, 2023). Key areas of novelty include offering more specific guidelines on when the model should defer to a human expert, and eliciting key information required for responsible assistance early in the conversation. We end by discussing the practical implications of our work, including where LLM policy prototyping can be situated in AI alignment pipelines, and approaches for scaling up policy prototyping efforts.
@@
  • A 15-week observational study with 9 mental health experts that surfaced opportunities for tight, collaborative feedback loops in LLM policy design.
@@
Figure 2. Research Process Overview. Our work proceeded in 4 phases: (1) a 15-week observational study with 9 mental health experts (19 workshops) led to (2) conceptualization of LLM policy prototyping. We then (3) designed and built PolicyPad and (4) evaluated it through 8 policy prototyping sessions with 22 experts (10 mental health, 12 legal).
Four sticky notes in a line with arrows connecting them from left to right. Each sticky note describes a phase of our work. Our work proceeded in 4 phases: (1) a 15-week observational study with 9 mental health experts (19 workshops) led to (2) conceptualization of LLM policy prototyping. We then (3) designed and built PolicyPad and (4) evaluated it through 8 policy prototyping sessions with 22 experts (10 mental health, 12 legal)
@@ 2. Related Work
2.1. Co-Designing AI Systems
As AI systems simultaneously broaden and deepen their impact across society, there is an increasingly clear and urgent need to seek input on AI development from beyond AI developers (Suresh et al., 2024; Delgado et al., 2023; Clegg, 2023; Feng et al., 2025; Huang et al., 2024; OpenAI, 2025a; Cheong et al., 2024; Feng et al., 2024). Co-design provides an appropriate set of methods and frameworks to draw upon for diverse participants to collaboratively shape AI systems. Sanders and Stappers describe co-design as “the creativity of designers and people not trained in design working together in the design development process” (Sanders and Stappers, 2008).
Co-design can be used for AI development and evaluation in many ways. For example, Lin et al. (Lin and Van Brummelen, 2021) hosted co-design workshops with K–12 teachers to identify promising opportunities to integrate AI into core curricula. Frontier model developers have co-designed safety benchmarks for chemical and biological risks with virologists and national security experts through special partnerships with organizations like Gryphon Scientific (Anthropic, 2025; Hurst et al., 2024). Overall, the ‘co’ in ‘co-design’ can refer to a variety of terms that capture the spirit of this approach, including collaborative, cooperative, collective, or connective (Zamenopoulos and Alexiou, 2018). In our work, we emphasize the collaborative nature of co-design by supporting collaboration between domain experts and HCI researchers, as well as collaboration amongst domain experts themselves.
Co-design is often discussed interchangeably with the related practice of participatory design (PD) (Zamenopoulos and Alexiou, 2018; Yu, 2025). While the two share similar aspirations and commitments, they have noteworthy differences in scope and procedure. Yu et al. (Yu, 2025) argue that co-design encompasses a broader set of methods shaped by contributions from domains like product design, service design, and healthcare, while PD retains a more political identity, with strong commitments to democratic legitimacy, power redistribution, and long-term relational engagement. Indeed, these commitments are reflected in work in Participatory AI (Kuo et al., 2024; Suresh et al., 2024; Delgado et al., 2023; Clegg, 2023; Barnett et al., 2025a; Ajmani et al., 2025; Sloane et al., 2022). Additionally, participatory design creates a hybrid space that is neither in the stakeholders’ nor facilitators’ domain to allow collaboration to happen on “even ground” for all (Delgado et al., 2023), whereas co-design aims to explicitly draw upon the expertise of collaborating stakeholders (Long et al., 2021; Madaio et al., 2020; Lin and Van Brummelen, 2021; Tseng et al., 2025). Finally, co-design often focuses on designing products or artifacts (Kleinsmann and Valkenburg, 2008), while participatory design may have a broader focus on sociotechnical systems that may contain those products or artifacts (Delgado et al., 2023).
Drawing from co-design, we work closely with domain experts in mental health and law to prototype LLM behavioral policies for those domains. We contribute to existing literature on co-designing AI systems by focusing co-design efforts around an artifact that is increasingly recognized as central to AI alignment and governance strategies (OpenAI, 2025b; Anthropic, 2023; He et al., 2025; Zhang et al., 2025b): LLM behavioral policies.
2.2. Scenarios as Decision-Making and Anticipatory Tools
Scholars across disciplines have found scenarios to be a valuable tool in their research and practice. Two complementary uses of scenarios are to 1) directly support decision making, and 2) help anticipate potential future outcomes that can then lead to more informed decisions.
Scenarios have been shown to be helpful in grounding human decision-making, case studies are scenario-based curricula commonly used to train practitioners to sharpen their decision-making abilities in fields such as law, medicine, and business (Molewijk et al., 2008; Farashahi and Tajeddin, 2018; Feng et al., 2023). In the legal domain specifically, cases represent past court decisions and are foundational for case law, in which courts rely on rulings in precedent cases to resolve new ones (Caputo, 2024; Kagan, 2019). In collaborative settings with more than one decision-maker, cases provide contextual details and nuances that serve as a medium for productive deliberation to establish common ground and reconcile differences (Ontanón and Plaza, 2006; Glez-Bedia et al., 2002). For example, PolicyCraft (Kuo et al., 2024) is an interactive system that uses case-based deliberation and voting to scaffold participatory policy design. Other applications of case-based deliberation include legal adjudication (Carpenter, 1917), medical diagnosis (Schaekermann et al., 2019), and content moderation (Fan and Zhang, 2020; Pan et al., 2022).
Scenarios are also useful for envisioning potential uses of technology and anticipating technology’s societal consequences. Scenario-based design and storyboarding are commonly used early in the human-centered design process as provocations for designers envision ways a technology can meet user needs (Figma, 2025; Carrol, 1999; Bodker, 1999; Hooper and Hsia, 1982; Kuo et al., 2023). Barnett et al. (Barnett et al., 2025b) surveyed scenario-building methods in computer science research over the past decade and identified five main ways computing researchers use scenarios: 1) to gather stakeholder needs and values, 2) to empower marginalized groups to imagine technology futures, 3) to provoke ethical reflection and promote critical awareness, 4) to anticipate threats and risks of these technologies, and 5) to explore perceptions and impacts of novel technologies before they launch. Interactive systems have been developed by the HCI community (e.g., Blip (Pang et al., 2024), Farsight (Wang et al., 2024)) to help developers anticipate impacts of their work by retrieving existing scenarios or generating new ones.
In our work, we primarily use scenarios to aid expert deliberation and decision-making on what appropriate model behavior should look like in high-stakes, domain-specific contexts. We develop our scenarios by sampling datasets of realistic cases in mental health and legal practice introduced by prior work (Lamparth et al., 2025; Cheong et al., 2024). Our work builds on prior work by enabling interactive scenarios that experts can extend, collaboratively analyze, and use to interrogate model behavior. As they do so and work with others to collaboratively design an LLM policy, experts may also become better equipped to anticipate consequences and risks of LLM use in their domains.
2.3. Human-Centered Approaches to AI Alignment and Safety
AI alignment broadly refers to efforts to align the behavior of AI systems to human preferences and values (Gabriel, 2020; Ngo et al., 2022). Relatedly, AI safety aims to reduce risks that arise from the development and deployment of AI systems (Kasirzadeh, 2024; Anderljung et al., 2023; Bengio et al., 2025). Researchers increasingly recognize the importance of human-centered approaches, grounded in users’ personal and collective needs, in both areas (Terry et al., 2024; Deng et al., 2025; Morris, 2025; Feng et al., 2024; Lam et al., 2024; Feng et al., 2025; Kuo et al., 2023).
A growing body of work responding to this recognition focuses on empowering non-AI experts to surface problematic AI behavior through evaluating, auditing, and red-teaming model outputs (Deng et al., 2025; Lam et al., 2022; DeVrio et al., 2022; Shen et al., 2021). For example, DeVrio et al. (DeVrio et al., 2022) investigated everyday users’ strategies for uncovering harmful algorithmic behavior and found that 1) these strategies were strongly influenced by personal experiences with societal bias, and 2) collaborative sensemaking between multiple users is a promising approach for user-driven audits. WeAudit (Deng et al., 2025) presents a workflow and platform for everyday users to collectively audit text-to-image models and generate an audit report with actionable insights to AI developers. Developers have also employed hybrid human-AI red-teaming campaigns (Anthropic, 2025; Google, 2025; Hurst et al., 2024; Zhang et al., 2025a) or purely automated ones (Perez et al., 2022; Zhou et al., 2025). Across all of these approaches, prompt engineering has been extensively used as a technique to probe model behavior (Kim et al., 2024; Arawjo et al., 2024; Ganguli et al., 2022).
Auditing and red-teaming at the output level is often a game of whack-a-mole: there is seemingly an endless stream of problematic behaviors to patch and new ones are constantly surfacing. Researchers have thus developed complementary approaches where they define higher-level behavioral policies—often called “model specs” (OpenAI, 2025b) or “constitutions” (Anthropic, 2023)—and train them into the model (Bai et al., 2022; Huang et al., 2024; OpenAI, 2025a; Guan et al., 2024). These policies can be a promising artifact model behavior governance because they 1) can be easily authored and understood by people (He et al., 2025; Zhang et al., 2025b), 2) can serve as guidelines for human red-teamers and data annotators to improve other parts of the alignment pipeline (OpenAI, 2025b), and 3) are shown to be effective in inducing more safe and aligned model behavior, even before the model is red-teamed (Guan et al., 2024; Bai et al., 2022). However, these policies are not without the limitations. Vague or poorly written policies can be misinterpreted by the model (He et al., 2025), or policies may contain conflicting statements that give the model noisy signals during training (Zhang et al., 2025b). As such, designing these policies should be an iterative and continual process (OpenAI, 2025a).
Our work focuses on iteratively improving the design of LLM behavioral policies by inviting experts to contribute to them directly. While our primary focus is policy design, we are also inspired by approaches taken by literature on end-user model auditing and red-teaming. Specifically, we design affordances for experts to collaboratively test, critique, and discuss model outputs in response to policy changes to help them envision and implement policy improvements.